OpenStudy (anonymous):

a circle has an area of 153.86 square meters. What is the radius. ( a = pi r squared) and what is the circumference ( 2 pi r)

OpenStudy (anonymous):

You're given the equations, just work backwards.

Directrix (directrix):

Area of a Circle = pi r^2 153.86 = pi r^2 49 = r^2 7 = r meters --- Circumference = 2 pi r C = 2 (3.14) (7) C = 43.96 meters